# Editing Across Languages: A Survey of Multilingual Knowledge Editing

### v1

@@ Abstract
While Knowledge Editing has been extensively studied in monolingual settings, it remains underexplored in multilingual contexts. This survey systematizes recent research on Multilingual Knowledge Editing (MKE), a growing subdomain of model editing focused on ensuring factual edits generalize reliably across languages. We present a comprehensive taxonomy of MKE methods, covering parameter-based, memory-based, fine-tuning, and hypernetwork approaches. We survey available benchmarks, summarize key findings on method effectiveness and transfer patterns, identify challenges in cross-lingual propagation, and highlight open problems related to language anisotropy, evaluation coverage, and edit scalability. Our analysis consolidates a rapidly evolving area and lays the groundwork for future progress in editable language-aware LLMs.
@@ 1 Introduction
While KE methods have made impressive progress in monolingual settings, they largely overlook a critical real-world dimension: multilingualism. Multilingual LLMs are increasingly deployed in cross-lingual applications, and it is insufficient for a knowledge update to take effect only in a single language. Ideally, an edit in one language should propagate across all supported languages, ensuring consistent and equitable access to updated knowledge regardless of linguistic background. This challenge has given rise to the emerging field of multilingual knowledge editing (MKE), which seeks to synchronize factual updates across languages while preserving unrelated knowledge. Although several recent works Xu et al. (2023); Wang et al. (2024c); Wu et al. (2025) have begun addressing this problem, research remains fragmented and unsynthesized.
In this paper, we present the first comprehensive survey and comparative study of methods and benchmarks in multilingual knowledge editing. We review recent advancements in (i) proposed MKE methodologies, including parameter editing methods, memory-based methods, fine-tuning methods, and hypernetwork-based methods; (ii) benchmark datasets designed to evaluate multilingual knowledge editing effectiveness and cross-lingual generalization; (iii) comparative empirical findings and evaluations across methods; and (iv) key open challenges and future directions toward building LLMs that are reliably editable across languages. Our goal is to provide a structured foundation for future research and help unify this fast-growing but currently scattered domain.
While several surveys have examined knowledge editing in LLMs, they focus primarily on monolingual settings or broad architectural taxonomies. Mazzia et al. (2024) provide a general overview of neural model editing across modalities. Wang et al. (2024b) emphasize technical classifications and optimization formulations, while Zhang et al. (2024) focus on monolingual editing and interpretability. None of these surveys consider the specific challenges of multilingual propagation or evaluation.
By contrast, our survey focuses explicitly on the multilingual dimension. We contribute (i) a multilingual reinterpretation of existing method families, (ii) a synthesis of empirical trends observed across MKE studies, and (iii) targeted visualizations comparing methods, languages, and model scales. Our work complements and extends prior surveys by highlighting a critical but underexplored axis of LLM editability.
@@ 2 Multilingual Knowledge Editing
Let =$\mathcal{L}=\{\ell_{1},\ell_{2},\ldots\}$ denote the set of languages, and represent a fact as a triple $(s,r,o_{\text{new}})$. Given queries $Q_{\ell_{s}}$ in a source language $\ell_{s}$, and corresponding queries $Q_{\ell_{t}}$ in target languages $\ell_{t}\in\mathcal{L}$, the goal is to find updated parameters $\theta^{\prime}$ [^1] Non-parametric methods apply edits at inference time through prompts or external memory without modifying $\theta$. such that:
@@ 3 Methods in Multilingual KE
Work in multilingual knowledge editing (MKE) builds on monolingual techniques and introduces new methods to support cross-lingual consistency. These approaches can be broadly grouped into four methodological families: (i) Parameter Editing, which modifies model weights directly; (ii) Memory-based, which use in-context learning or external memory without updating parameters; (iii) Fine-tuning, including full and instruction-tuned updates; and (iv) Hypernetwork-based, which generate edits via auxiliary networks. This taxonomy supports a structured comparison of design choices and trade-offs in MKE.
@@ 3.1.1 General KE Methods Adapted for MKE
ROME Meng et al. (2023a) introduced the locate-then-edit paradigm, identifying key activation pathways and applying a low-rank update to model weights. MEMIT Meng et al. (2023b) extended ROME to support batch editing of multiple facts using rank-one updates $\{\Delta\theta_{k}\}_{k=1}^{K}$ distributed across layers. While it preserves locality and improves efficiency, its effectiveness across languages is limited by language-specific neuron activations. Similarly, KnowledgeNeuron (KN) Dai et al. (2022) and PMET Li et al. (2024) followed the locate-and-edit paradigm by optimizing $\Delta\theta$ to encode factual edits. Wang et al. (2024a); Zhang et al. (2025) adapted ROME, MEMIT, and PMET to multilingual models. While these adaptations preserve edit locality and maintain computational efficiency, they suffer from poor cross-lingual propagation due to language-specific activation patterns and representational anisotropy.
@@ 3.1.3 Summary
Strengths Parameter editing methods offer localized updates that reduce interference with unrelated knowledge and are efficient for factual edits without full retraining. They are also compatible with standard transformer architectures, making them easy to integrate into multilingual LLM pipelines.
Weaknesses These methods rely on shared internal representations across languages, which are often weak or inconsistent. As a result, they do not guarantee reliable cross-lingual propagation; edits made in one language may fail to transfer to others. Their performance is also sensitive to tokenization and neuron alignment, particularly in morphologically rich or low-resource languages. While MKE-specific approaches (e.g LU-LAFNs) address these issues through attention or neuron-level interventions, key theoretical challenges remain.
@@ 3.2 Memory-based Methods
3.2.1 In-Context Editing (IKE and ICE)
In-context editing (ICE) Zheng et al. (2023) modifies model behavior by injecting factual corrections into the input prompt, without changing model weights ($\theta^{\prime}=\theta$). Zero-shot ICE uses single factual statements, while IKE extends this with few-shot demonstrations. These methods can generalize across languages if equivalent demonstrations are available in each language.  Nie et al. (2025) for example applied in-context editing in 53 languages.
@@ 3.2.2 Memory-Retriever Editing
Memory-Retriever methods attach an external memory to the LLM and retrieve language-specific facts $\{m_{i}\}$ at inference time, producing $f(q,\{m_{i}\};\theta)\approx o_{\text{new}}$ without modifying weights. This design supports multilingual edits by isolating updates outside the model, but effectiveness depends on retriever quality. Variants differ in how memory is used: SERAC Mitchell et al. (2022b) overrides predictions with matched entries. ReMaKE Wang et al. (2024c) uses retrieved memories to guide lightweight updates. MQA-KEAL Ali et al. (2025) performs multi-hop reasoning over memory chains. Cross-lingual retrievers like Mello-CL Khandelwal et al. (2024) boost recall in multilingual and low-resource contexts.
@@ 3.2.3 Summary
Strengths Memory-based methods avoid parameter updates, minimizing interference with unrelated cross-lingual knowledge. Their edits are applied externally at inference time, enabling high flexibility for dynamic and reversible updates. Because corrections are decoupled from internal weights, these approaches can be applied independently to any language, making them naturally suitable for multilingual deployments.
Weaknesses Their effectiveness depends on language-specific components: prompt understanding and instruction following for in-context methods, or retriever accuracy and memory quality for retrieval-based approaches. Neither guarantees cross-lingual propagation; corrections must be explicitly provided or retrieved in each target language. In addition, memory-retriever methods incur storage and latency costs as edits accumulate, while in-context editing requires constructing effective prompts across languages.
@@ 3.3.2 Instruction-Tuned Knowledge Editing
X-KDE Wu et al. (2025) extends LTE to multilingual settings via a two-stage framework: Cross-lingual Edition Instruction Tuning (XE-IT) trains on parallel data to transfer edits across languages, and Target-Language Preference Optimization (TL-PO) aligns outputs for consistency.
@@
Instruction-tuned approaches are the first systematic attempt to model editing as a generalizable capability. LTE is effective in monolingual contexts, while X-KDE introduces a supervised framework for cross-lingual propagation and alignment.
@@ 3.3.3 Summary
Strengths Fine-tuning methods are widely used for knowledge editing due to their simplicity and strong factual reliability. They are compatible with pre-trained transformer architectures and require minimal structural changes. LoRA-based fine-tuning improves efficiency by updating a small parameter subset. Instruction-tuned approaches, such as LTE and X-KDE, treat editing as a learnable task, offering more control over behavior and reducing unintended side effects. Notably, X-KDE is the first supervised fine-tuning framework specifically designed to promote cross-lingual consistency.
Weaknesses Fine-tuning provides no guarantees for cross-lingual propagation; generalization to target languages depends on implicit internal alignment. Large parameter updates also risk altering unrelated knowledge. While instruction-tuned methods mitigate this by explicitly modeling the editing process, they rely on large, curated datasets and careful instruction design.
@@ 3.4 Hypernetwork-Based Methods
Hypernetwork-based methods edit knowledge by training auxiliary networks to generate updates for a model. Instead of manual weight changes, the hypernetwork learns to produce edits from query-output pairs. MEND Mitchell et al. (2022a) introduced this by training a hypernetwork $h_{\phi}$ to produce updates $\Delta\theta=h_{\phi}(q,o_{\text{new}})$ such that $f(q;\theta+\Delta\theta)\approx o_{\text{new}}$, while minimizing interference. KnowledgeEditor De Cao et al. (2021) added regularization to stabilize edits and improve generalization. LiME Xu et al. (2023) adapted this paradigm for multilingual editing by conditioning the hypernetwork on source language $\ell_{s}$, generating edits in language-specific subspaces: $f(q;\theta+h_{\phi}(q,o_{\text{new}},\ell_{s}))\approx o_{\text{new}}$. By modeling language anisotropy explicitly, LiME improved cross-lingual stability.
@@ 3.4.1 Summary
Strengths Hypernetwork-based methods enable meta-learning of edits without manual parameter selection, offering fast, localized updates with minimal disruption. Language-conditioned variants such as LiME introduce explicit mechanisms to improve edit stability across languages.
Weaknesses The effectiveness of hypernetwork-based methods depends heavily on the quality and coverage of training data. While MEND and KnowledgeEditor were designed for monolingual models and used as multilingual baselines, Beniwal et al. (2024) showed their limited cross-lingual transfer. LiME mitigates this by conditioning on the source language, but its reliance on language-specific subspaces may limit generalization across distant language pairs.
@@ 3.5 Summary of Methods
Table 1: Comparison of multilingual knowledge editing methods across core evaluation criteria. Edit Rel: edit reliability in source language; X-ling Prop: cross-lingual propagation; Locality: preservation of unrelated knowledge; Param. Mod: modifies model weights; Dyn. Flex: supports inference-time edit control. ✓= strong; $\sim$ = partial; ✗= absent
@@
MzSRE Wang et al. (2024c) 12 Translated via Google; used for all four metrics
@@
CROLIN-MQuAKE Khandelwal et al. (2024) 7 Translations verified via BLEU and human checks
@@
Eva-KELLM Wu et al. (2023) 2 ? Generated from Counterfact using ChatGPT
@@
CKnowEdit Fang et al. (2024) 1 Chinese-focused cultural and linguistic error correction
Table 2: Multilingual knowledge editing benchmarks. Langs. = number of languages. Human? = human annotation or translation verification
@@ 5.1 Comparative Evaluation
Parameter-based methods such as ROME and MEMIT perform strongly in monolingual settings but struggle in multilingual contexts, with frequent knowledge conflicts and degraded cross-lingual transfer Wang et al. (2024c); Wu et al. (2025); Zhang et al. (2025); Wei et al. (2025); Xie et al. (2024). LU-LAFNs achieves perfect locality and strong reliability by identifying and modifying shared factual neurons across languages. A key reason for its superior locality is how updates are applied: after optimization, the updates are stored externally and only injected at inference time if a matching subject is detected in the user prompt. This minimizes the risk of unintended changes to unrelated knowledge. However, like other parameter editing methods, LU-LAFNs shows weaker cross-lingual generalization and portability.
@@
Memory-based methods like IKE and ReMaKE avoid parameter updates and instead rely on retrieval or prompting. ReMaKE achieves excellent locality but suffers from low reliability in many multilingual scenarios, with performance often dropping below 50%. In contrast, IKE demonstrates relatively poor locality, as it heavily depends on the quality and selection of few-shot exemplars Wang et al. (2024a, c); Wu et al. (2025). Additionally, the applicability of ReMaKE across model families remains limited, its performance degrades significantly when applied to models such as BLOOM Zhang et al. (2025), indicating strong dependence on the underlying model’s capabilities. Despite their limitations, prompting-based approaches remain surprisingly competitive: Nie et al. (2025) demonstrated that simple in-context editing achieves reasonable cross-lingual edit success across 53 languages, even without model weight modifications.
@@
Figure 2: The figure shows average Reliability and Generality for MEMLA Xie et al. (2024) across twelve languages, illustrating a typical directionality pattern: edits in high-resource or related languages transfer better than those in low-resource or distant ones.
5.2 Cross Lingual Transferability

### v2

@@ Abstract
While Knowledge Editing has been extensively studied in monolingual settings, it remains underexplored in multilingual contexts. This survey systematizes recent research on Multilingual Knowledge Editing (MKE), a growing subdomain of model editing focused on ensuring factual edits generalize reliably across languages. We present a comprehensive taxonomy of MKE methods, covering parameter-based, memory-based, fine-tuning, and hypernetwork approaches. We survey available benchmarks, summarize key findings on method effectiveness and transfer patterns, and identify persistent challenges such as cross-lingual propagation, language anisotropy, and limited evaluation for low-resource and culturally specific languages. We also discuss broader concerns such as stability and scalability of multilingual edits. Our analysis consolidates a rapidly evolving area and lays the groundwork for future progress in editable language-aware LLMs.
@@ 1 Introduction
While KE methods have made impressive progress in monolingual settings, they overlook a critical real-world dimension: multilinguality. Multilingual LLMs are increasingly deployed in cross-lingual applications, and it is insufficient for a knowledge update to take effect only in a single language. Ideally, an edit in one language should propagate across all supported languages, ensuring consistent and equitable access to updated knowledge regardless of linguistic background. This challenge has given rise to the emerging field of Multilingual Knowledge Editing (MKE), which seeks to synchronize factual updates across languages while preserving unrelated knowledge.
In this paper, we present the first comprehensive survey of methods and benchmarks for multilingual knowledge editing. We review recent advancements in (i) proposed MKE methodologies, including parameter editing methods, memory-based methods, fine-tuning methods, and hypernetwork-based methods; (ii) benchmark datasets designed to evaluate multilingual knowledge editing effectiveness and cross-lingual generalization; (iii) empirical findings on method effectiveness and cross-lingual generalization; and (iv) key open challenges and future directions toward building LLMs that are reliably editable across languages. Our goal is to provide a structured foundation for future research and help unify this fast-growing but currently scattered domain.
While several surveys have examined knowledge editing in LLMs, they focus primarily on monolingual settings or broad architectural taxonomies. Mazzia et al. (2024) provide a general overview of neural model editing across modalities. Wang et al. (2024c) emphasize technical classifications and optimization formulations, while Zhang et al. (2024a) focus on monolingual editing and interpretability. None of these surveys consider the specific challenges of multilingual propagation or evaluation.
By contrast, our survey focuses explicitly on the multilingual dimension. We contribute (i) a multilingual reinterpretation of existing method families, (ii) a synthesis of empirical trends observed across MKE studies, and (iii) targeted visualizations comparing methods, languages, and model scales. Our work complements and extends prior surveys by highlighting multilingual knowledge editing as a critical but underexplored area.
@@ 2 Multilingual Knowledge Editing
Let $\mathcal{L}=\{\ell_{1},\ell_{2},\ldots\}$ denote the set of languages, and represent a fact as a triple $(s,r,o_{\text{new}})$. Given queries $Q_{\ell_{s}}$ in a source language $\ell_{s}$, and corresponding queries $Q_{\ell_{t}}$ in target languages $\ell_{t}\in\mathcal{L}$, the goal is to find updated parameters $\theta^{\prime}$ [^1] Non-parametric methods apply edits at inference time through prompts or external memory without modifying $\theta$. such that:
@@ 3 Methods in Multilingual KE
Work in multilingual knowledge editing (MKE) builds on monolingual techniques and introduces new methods to support cross-lingual consistency. These approaches can be broadly grouped into four methodological families: (i) Parameter Editing, which modifies model weights directly; (ii) Memory-based methods, which use in-context learning or external memory without updating parameters; (iii) Fine-tuning, including full and instruction-tuned updates; and (iv) Hypernetwork-based, which generate edits via auxiliary networks. This taxonomy supports a structured comparison of design choices and trade-offs in MKE.
@@ 3.1.1 General KE Methods Adapted for MKE
ROME Meng et al. (2023a) introduced the locate-then-edit paradigm, identifying key activation pathways and applying a low-rank update to model weights. MEMIT Meng et al. (2023b) extended ROME to support batch editing of multiple facts using rank-one updates $\{\Delta\theta_{k}\}_{k=1}^{K}$ distributed across layers. While it preserves locality and improves efficiency, its effectiveness across languages is limited by language-specific neuron activations. Similarly, KnowledgeNeuron Dai et al. (2022) and PMET Li et al. (2024) followed the locate-and-edit paradigm by optimizing $\Delta\theta$ to encode factual edits. Wang et al. (2024a); Zhang et al. (2025) adapted ROME, MEMIT, and PMET to multilingual models. While these adaptations preserve edit locality and maintain computational efficiency, they suffer from poor cross-lingual propagation due to language-specific activation patterns and representational anisotropy.
@@ 3.1.3 Summary
Strengths: Parameter editing methods offer localized updates that reduce interference with unrelated knowledge and are efficient for factual edits without full retraining. They are also compatible with standard transformer architectures, making them easy to integrate into multilingual LLM pipelines.
Weaknesses: These methods rely on shared internal representations across languages, which are often weak or inconsistent. As a result, they do not guarantee reliable cross-lingual propagation; edits made in one language may fail to transfer to others. Their performance is also sensitive to tokenization and neuron alignment, particularly in morphologically rich or low-resource languages. While MKE-specific approaches (e.g LU-LAFNs) address these issues through attention or neuron-level interventions, key theoretical challenges remain.
@@ 3.2 Memory-based Methods
3.2.1 In-Context Editing (ICE & IKE)
In-context editing Zheng et al. (2023) modifies model behavior by injecting factual corrections into the input prompt, without changing model weights. Zero-shot ICE uses single factual statements, while IKE extends this with few-shot demonstrations. These methods can generalize across languages if equivalent demonstrations are available in each language.  Nie et al. (2025) for example applied in-context editing in 53 languages.
However, these methods are constrained by context length: edits must be carried in the prompt, which reduces space for task inputs and outputs. This issue is especially severe in languages with high tokenizer fertility, where facts occupy more tokens, and in sequential editing scenarios, where multiple edits quickly accumulate.
@@ 3.2.2 Memory-Retriever Editing
Memory-Retriever methods attach an external memory to the LLM and retrieve language-specific facts $\{m_{i}\}$ at inference time, producing $f(q,\{m_{i}\};\theta)\approx o_{\text{new}}$ without modifying weights. This design supports multilingual edits by isolating updates outside the model, but effectiveness depends on retriever quality. Variants differ in how memory is used: SERAC Mitchell et al. (2022b) overrides predictions with matched entries. ReMaKE Wang et al. (2024d) uses retrieved memories to guide lightweight updates. MQA-KEAL Ali et al. (2025) performs multi-hop reasoning over memory chains. Cross-lingual retrievers like Mello-CL Khandelwal et al. (2024) boost recall in multilingual and low-resource contexts. [^2] Other approaches such as GRACE Hartvigsen et al. (2023) and WISE Wang et al. (2024b) also support handling multiple edits while maintaining locality. We did not include them in our main taxonomy since they were developed and evaluated only in English and do not address cross-lingual generalization, but they represent promising directions for future MKE research.
@@ 3.2.3 Summary
Strengths: Memory-based methods avoid parameter updates, minimizing interference with unrelated cross-lingual knowledge. Their edits are applied externally at inference time, enabling high flexibility for dynamic and reversible updates. Because corrections are decoupled from internal weights, these approaches can be applied independently to any language, making them naturally suitable for multilingual deployments.
Weaknesses: Their effectiveness depends on language-specific components: prompt understanding and instruction following for in-context methods, or retriever accuracy and memory quality for retrieval-based approaches. Neither guarantees cross-lingual propagation; corrections must be explicitly provided or retrieved in each target language. In addition, memory-retriever methods incur storage and latency costs as edits accumulate, while in-context editing requires constructing effective prompts across languages.
@@ 3.3.2 Instruction-Tuned Knowledge Editing
X-KDE Wu et al. (2025) extended LTE to multilingual settings via a two-stage framework: Cross-lingual Edition Instruction Tuning (XE-IT) trains on parallel data to transfer edits across languages, and Target-Language Preference Optimization (TL-PO) aligns outputs for consistency.
@@
Instruction-tuned approaches are the first systematic attempt to model editing as a generalizable capability. LTE is effective in monolingual contexts, while X-KDE introduced a supervised framework for cross-lingual propagation and alignment.
@@ 3.3.3 Summary
Strengths: Fine-tuning methods are widely used for knowledge editing due to their simplicity and strong factual reliability. They are compatible with pre-trained transformer architectures and require minimal structural changes. LoRA-based fine-tuning improves efficiency by updating a small parameter subset. Instruction-tuned approaches, such as LTE and X-KDE, treat editing as a learnable task, offering more control over behavior and reducing unintended side effects. Notably, X-KDE is the first supervised fine-tuning framework specifically designed to promote cross-lingual consistency.
Weaknesses: Fine-tuning provides no guarantees for cross-lingual propagation; generalization to target languages depends on implicit internal alignment. Large parameter updates also risk altering unrelated knowledge. While instruction-tuned methods mitigate this by explicitly modeling the editing process, they rely on large, curated datasets and careful instruction design.
@@ 3.4 Hypernetwork-Based Methods
Hypernetwork-based methods edit knowledge by training auxiliary networks to generate updates. Instead of manual weight changes, hypernetwork learns to produce edits from query-output pairs. MEND Mitchell et al. (2022a) introduced this by training a hypernetwork $h_{\phi}$ to produce updates $\Delta\theta=h_{\phi}(q,o_{\text{new}})$ such that $f(q;\theta+\Delta\theta)\approx o_{\text{new}}$, while minimizing interference. KnowledgeEditor De Cao et al. (2021) added regularization to stabilize edits and improve generalization. LiME Xu et al. (2023) adapted this paradigm for multilingual editing by conditioning the hypernetwork on source language $\ell_{s}$, generating edits in language-specific subspaces: $f(q;\theta+h_{\phi}(q,o_{\text{new}},\ell_{s}))\approx o_{\text{new}}$. By modeling language anisotropy explicitly, LiME improved cross-lingual stability.
@@ 3.4.1 Summary
Strengths: Hypernetwork-based methods enable meta-learning of edits without manual parameter selection, offering fast, localized updates with minimal disruption. Language-conditioned variants such as LiME introduce explicit mechanisms to improve edit stability across languages.
Weaknesses: The effectiveness of hypernetwork-based methods depends heavily on the quality and coverage of training data. While MEND and KnowledgeEditor were designed for monolingual models and used as multilingual baselines, Beniwal et al. (2024) showed their limited cross-lingual transfer. LiME mitigates this by conditioning on the source language, but its reliance on language-specific subspaces may limit generalization across distant language pairs.
@@ 3.5 Summary of Methods
Table 1: Comparison of multilingual knowledge editing methods across core evaluation criteria. Edit Rel: edit reliability in source language; X-ling Prop: cross-lingual propagation; Locality: preservation of unrelated knowledge; Param. Mod: modifies model weights; Dyn. Flex: supports inference-time edit control. ✓= strong; $\sim$ = partial; ✗= absent.
@@
MzSRE Wang et al. (2024d) 12 Translated via Google; used for all four metrics
@@
MQuAKE Khandelwal et al. (2024) 7 Translations verified via BLEU and human checks
@@
DocTer Wu et al. (2023) 2 ? Generated from Counterfact using ChatGPT
@@
CKnowEdit Fang et al. (2024) 1 Chinese cultural and linguistic error correction
Table 2: Multilingual knowledge editing benchmarks. Langs. = number of languages. Human? = human annotation or translation verification.
@@ 5.1 Comparative Evaluation
Parameter-based methods such as ROME and MEMIT perform strongly in monolingual settings but struggle in multilingual contexts, often causing interference and degraded cross-lingual transfer Wang et al. (2024d); Wu et al. (2025); Zhang et al. (2025); Wei et al. (2025); Xie et al. (2024). LU-LAFNs achieves perfect locality and strong reliability by identifying and modifying shared factual neurons across languages. A key reason for its superior locality is how updates are applied: after optimization, the updates are stored externally and only injected at inference time if a matching subject is detected in the user prompt. This minimizes the risk of unintended changes to unrelated knowledge. However, like other parameter editing methods, LU-LAFNs shows weaker cross-lingual generalization and portability.
@@
Memory-based methods like IKE and ReMaKE avoid parameter updates and instead rely on retrieval or prompting. ReMaKE achieves excellent locality but suffers from low reliability in many multilingual scenarios, with performance often dropping below 50%. In contrast, IKE demonstrates relatively poor locality, as it heavily depends on the quality and selection of few-shot exemplars Wang et al. (2024a, d); Wu et al. (2025). Additionally, the applicability of ReMaKE across model families remains limited, its performance degrades significantly when applied to models such as BLOOM Zhang et al. (2025), indicating strong dependence on the underlying model’s capabilities. Despite their limitations, prompting-based approaches remain surprisingly competitive: Nie et al. (2025) demonstrated that simple in-context editing achieves reasonable cross-lingual edit success across 53 languages, even without model weight modifications.
@@
Figure 2: Average Reliability and Generality for MEMLA Xie et al. (2024) across twelve languages, illustrating a typical directionality pattern: edits in high-resource or related languages transfer better than those in low-resource or distant ones.
5.2 Cross-Lingual Transferability